• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 01:00:23 on July23,2003

    Entry number 106297

    Followups:

    keyword=Cell has probably fail

    12:25: The RTds have been reading wrong values (0 degree, 989 degree)... So I turned off the 3 transverse lasers.
    At the same time, the septum began to warm up and the rate was much higher that it used to be (about 45 kHz instead of 8 kHz for I=2.5 uA).
    Beam was on and centered. Raster was on, according to the scope (but we did not do a spot++ check so there is a slim possibility that something went bad in the optics during the previous hour down and the real spot sized changed). These symptoms appers at 12:25, about 5 min after the beam was restored (with a proper ramp up).
    The ref cell seems to hold vacuum.
    We called for an access. We also called radcon.
    We called Jian-Ping who suggested to see with one transverse laser if we have absorption but no deep appeared in the spectrum. We are now waiting for the control access.